### v1

# LangGPT: Rethinking Structured Reusable Prompt Design Framework for LLMs from the Programming Language

Abstract
LLMs have demonstrated commendable performance across diverse domains. Nevertheless, formulating high-quality prompts to effectively instruct LLMs poses a challenge for non-AI experts. Existing research in prompt engineering suggests somewhat fragmented optimization principles and designs empirically dependent prompt optimizers. Unfortunately, these endeavors lack a structured design template, incurring high learning costs and resulting in low reusability. Inspired by structured reusable programming languages, we propose LangGPT, a dual-layer prompt design framework as the programming language for LLMs. LangGPT has an easy-to-learn normative structure and provides an extended structure for migration and reuse. Experiments illustrate that LangGPT significantly enhances the capacity of LLMs to produce responses of superior quality compared to baselines. Moreover, LangGPT has proven effective in guiding LLMs to generate high-quality prompts. We have built a community on LangGPT to facilitate the tuition and sharing of prompt design. We also analyzed the ease of use and reusability of LangGPT through a community user survey.
1 Introduction
Large language models (LLMs) like ChatGPT Achiam et al. (2023); Schulman et al. (2023) can execute diverse tasks Sun et al. (2023b, c); Yu et al. (2023) based on powerful language comprehension, reasoning, and generation capabilities. Injecting domain knowledge also enables LLMs to perform domain-related specific tasks Wang et al. (2023a); Li et al. (2023b); Zhang et al. (2023); Ren et al. (2023). Fully unleashing these capabilities of LLMs requires premium quality prompts Eric (2022); Chen et al. (2023); Gajula (2023). Therefore, prompt engineering has attracted the attention of many researchers Varshney and Surla (2023); Meskó (2023); Wang (2023).
Prompt engineering is a typically empirical science, which mainly involves designing and optimizing the prompts given to LLMs. As LLMs have the capability of natural language understanding, it is possible to ask them to execute tasks through unstructured natural language instructions directly. On this basis, some tricks for prompt optimization were first explored by researchers. Bsharat et al. (2023) introduce 26 guiding principles designed to make LLMs perform better. In addition to these directly usable prompt optimization tricks, some researchers have also focused on optimizing prompts based on historical data. Sun et al. (2023a) guide LLMs to derive new prompts for a given instance from the incorrect reasoning, and then summarise the corresponding prompts for each instance as a reference for optimizing the original prompts. Pryzant et al. (2023) leverage mini-batches of data to form natural language “gradients” and utilize beam search and bandit selection procedure to edit the current prompt in the opposite semantic direction of the gradient. Fan et al. (2023) analyze a large prompt database and present an automatic prompt optimization framework.
Direct prompt optimization principles and methods based on historical data require a wealth of experience. Thus, these usually perform well only for specific tasks or domains. To improve generalization, some researchers have proposed adaptive prompt optimization methods. Guo et al. (2023) connect LLMs with evolutionary algorithms and proposes a novel framework for discrete prompt optimization, called EvoPrompt. Li et al. (2023a) design a multi-round dialogue alignment strategy and utilizes GPT-4 Achiam et al. (2023) to generation a readability prompt set. Meanwhile, they propose an efficient prompt screening metric that can filtrate high-quality prompts with linear complexity. Wang et al. (2023b) introduce PromptAgent which can reflect on model errors and generate constructive error feedback to induce precise expert-level insights and in-depth instructions. Hao et al. (2022) and Cheng et al. (2023) optimize prompts from the perspective of aligning human and LLMs’ preference styles.
Prompt optimization can significantly improve the performance of LLMs, but the reusability of quality prompts is poor due to the lack of systematic design. Consequently, some researchers have devised rules for the construction of prompts. Nigh (2023) collects a large number of quality prompts and summaries the CRISPE rule for prompt design. Zamfirescu-Pereira et al. (2023) take a prototype LLM-based chatbot design tool as the design probe, supporting non-AI-experts engage in “end-user prompt engineering”. Some researchers have designed prompt construction rules for applications in different domains. Cao et al. (2023) present various prompt templates on deep learning program repair tasks for ChatGPT. Yeh et al. (2022) reformulate the biomedical relation extraction task as a cloze-test task under a simple prompt formulation to systematically generate comprehensive prompts. Liu and Chilton (2022) have evaluated 5493 generations covering 51 themes and 51 styles throughout five experiments in the text-to-image task and summaries prompt design guidelines.
These methods, which are based on a great deal of experience in use, are mainly based on listed design rules and lack systematicity. In addition, these rules have strong domain relevance and model relevance, with low generalisability, flexibility, and reusability. To further unleash the performance of LLMs, some researchers have defined agents. Agents empower LLMs to use tools, acquire domain knowledge, retain long-term or short-term memories, and plan Xu et al. (2023); Xi et al. (2023); Park et al. (2023). Although agent methods Chase (2022); Hong et al. (2023); Wu et al. (2023) have systematically designed the key components of the prompt and reserved flexible customization interfaces, the learning costs are very high. In addition, it is difficult for non-AI experts to modify agent designs, so excellent designs are less reusable.
To promote LLM-based applications and further stimulate the potential of LLMs, we would like to design a high-quality, reusable prompt template. This template should be extensible and generalizable. In addition, the template ought to be easy to learn and easy to work with. Inspired by the belief that prompt is the programming language of the LLM era Alouani (2023); Mund (2023), we have designed Language for GPT-like LLMs (LangGPT), a prompt design framework as the programming language for LLMs. LangGPT refers to the systematic, prescriptive, and reusable properties of programming languages, and retains the flexibility and extensibility of natural languages. We analyzed the differences between natural languages and programming languages to determine the properties that the prompt template should have. LangGPT is designed as a dual-layer structure, consisting of modules and internal elements. Modules in LangGPT can be divided into two categories: inherent modules and extension modules. For inherent modules, we design the necessary internal elements of each module in detail and give example templates. In addition, for the extension modules, we unified the design of the basic internal elements. Experiments have demonstrated that LangGPT is better than baseline prompts for bootstrapping LLMs. Furthermore, LLMs can leverage this framework to automatically generate prompts as if generating codes.
In summary, the contributions of this work include:
  • We proposed a dual-layer structured prompt design framework LangGPT to improve the generalization and reusability of prompts. In addition, we wrote detailed rules for designing prompts based on LangGPT, reducing the learning cost of prompt design and enabling LLMs to generate high-quality prompts automatically.
  • We proved experimentally that the prompts designed based on LangGPT can better guide LLMs to execute tasks. Meanwhile, we exemplified that LangGPT can help LLMs generate high-quality prompts.
  • We have built an online community based on LangGPT, which provides usage documentation and prompt design interfaces. In addition, collecting and sharing excellent prompt cases promotes the exchange of LLM applications. We conducted a user experience survey in the community to verify the ease of use and reusability of LangGPT.
2 Prompt Design Rules with Programming Language
Programming languages are more standardized and reusable compared to natural languages. To design high-quality reusable prompts, we analyzed the differences between natural languages and programming languages and proposed prompt design rules.
2.1 Difference between Programming Languages and Natural Languages
While natural languages are primarily used for communication, programming languages are designed to define instructions for machines to execute tasks GeeksforGeeks (2023). The different purposes of application have led to very different contexts for the creation and evolution of the two languages.
  • 1.
    Target audience. Natural languages are spoken by humans to humans Grune et al. (2012), while programming languages are prepared by humans for machines Chakray (2018).
  • 2.
    Structure. The fact that computers can only understand fixed instructions requires programming languages to have strict, rigorous syntax and semantics. In opposition, natural languages have loose and flexible syntax and semantics, permitting creativity and variation and possessing a high degree of fault tolerance.
  • 3.
    Ambiguity. Natural languages are more ambiguous, but humans have the ability to clarify the meaning of expressions. For example, problems such as jumbled order and typos have less of an impact on reading, and mispronunciations can be understood Chakray (2018); Aho (2007). Programming languages are less ambiguous because they need to provide computers with clear instructions GeeksforGeeks (2023).
  • 4.
    Evolution and updating. Natural languages evolve naturally over time through human use and communication Sipser (1996). Natural languages are flexible in adding new words and meanings and discarding outdated usages Fromkin et al. (2018). Programming languages are specifically designed to communicate with machines Sebesta (2012). New syntax rules and features require explicit upgrades or releases Pratt et al. (1984).
Figure 1: Analogy between programming language and natural language prompt. The analogy between the two types of languages was analyzed in terms of their hierarchical structure. Circles of different sizes indicate different layers. Smaller circles indicate closer to the inner layers, corresponding to darker colors.
In summary, the main difference between the two languages is that natural languages are more vague and flexible, while programming languages are more standardized and precise. LLMs essentially perform a large number of computations and share many similarities with machines. Therefore, we propose LangGPT, a natural language programming framework for LLMs, by drawing on the characteristics of programming languages and combining the advantages of natural languages.
2.2 Rules for Prompt Design
We improve prompts by referring to the design ideas of programming languages. After analyzing the differences between programming languages and natural languages, we propose the design rules for prompts: (1) Prompts should have a regularised format. Flexible and ambiguous natural languages are difficult to understand for LLMs. Format-constrained prompts make users’ purpose and requirements more salient. (2) The structure of prompts should be extensible. Custom structures facilitate users to design suitable prompts according to their domains and tasks. (3) Specific requirements must be clear and complete. Both instructions and additional requirements should be explicit and complete to avoid misunderstandings or biases. (4) Languages should be flexible. Where the requirements are clear, flexible languages can be better adapted to different domains. In addition, flexible languages are easy for to learn users.
Scenario Prof. Cons. Goal Init. Ex. Wkflo. Skill Sug. Bkgrd. Style Outf.
Writing
Role-playing
Entertainment
Supplementary Learning
Prompt Optimisation
Prompt Hacking
Drawing
Business Operation
Table 1: Status of inherent module definitions. The table lists the eight categories of application scenarios we have defined so far and the modules defined for these scenarios. A ✓indicates that a corresponding module has been designed for this scenario. In contrast, a ✗ indicates that it was not designed.
3 Natural Language Programming Framework of Prompts for LLMs
Based on the design rules, we present LangGPT, the natural language programming framework with a dual-layer architecture for LLMs.
3.1 Overall Dual-layer Structure
To systematically design prompts that meet the rules, we have made full reference to the design ideas and structures of object-oriented programming languages Rentsch (1982); Lutz (2010). We consider the prompt as a software project and analogize the prompt design process with the software development process. The correspondence is shown in Figure 1.
Based on analogical analyses, it can be found that natural language prompts have a similar multi-level structure as programming languages. Therefore, we refer to the structure of programming languages and propose a dual-layer structure for prompt design, and define the notion of module and element for prompts.
A complete prompt contains several modules. Modules are similar to classes in programming languages, and each module represents an aspect of the requirements for LLMs. For instance, prompts can be augmented in terms of constraints, goals, profiles, etc. Within a module, a number of internal elements are included. Elements are similar to functions and properties in programming languages and represent the content of direct and specific instructions to LLMs. For example, “Output should be no more than 500 words” could be an element in a prompt that belongs to the module constraint.
A dual-layer structure can be a good way to standardize the formatting of prompts. However, the flexibility of natural languages would be lost if prompts are too strictly required to follow predefined standard modules and internal elements. In addition, it will reduce the generalisability of LangGPT to different tasks in different domains, which is not conducive to the reuse of quality prompts. To solve these problems, we divided the types of modules and elements into prompts. We defined inherent module and basic element as the predefined dual-layer prompt template. In addition, we constructed extension module and custom element that support customization. We provide both Markdown Gruber (2012) and JSON Pezoa et al. (2016) formats for inherent modules and extension modules. Furthermore, we have written basic elements for different modules and defined rules for writing custom elements.
3.2 Tectonics of Inherent Modules
The module serves as a connection between the complete prompt and the instruction unit and plays a very important role in controlling the structure of the prompt.
We define inherent modules for critical aspects that are required for almost all prompts. Furthermore, we define inherent modules for certain scenarios that are relevant to the application for ease of learning and use. Table 1 shows the inherent modules we defined for some scenarios.
In Table 1, Prof. indicates what is expected of LLMs in terms of roles, including profiles, character portraits, etc. Cons. denotes constraints or attention, i.e., ranges that LLMs are not allowed to exceed and requirements that must be met when generating responses, etc. Goal lists the goals that the user wants to achieve, which is what the LLMs need to accomplish. Init. is called initialization to inform LLMs that they are about to start a dialogue. Sometimes a specified first sentence is also given in this module. Ex. gives LLMs input-output pairs as examples to learn from. Wkflo. instructs the workflow when executing a task, similar to the CoT approach Wei et al. (2023). It is often necessary to instantiate this module when the task requirements are more complex. Skill is used to suggest to LLMs the skills they possess. LLMs that have undergone tool learning can be guided to invoke tools to execute tasks more accurately. In addition, we plan to provide the ability to use tools under this module in future work, with reference to the design of agent tools Chase (2022); Hong et al. (2023). Sug. includes suggestions and behavioral planning for LLMs. This module focuses on listing common scenarios and giving behaviors or responses that LLMs can take in such situations. Bkgrd. indicates the background information and the memories that LLMs are required to have when performing their tasks. Style qualifies the style of responses generated by LLMs. Outf. defines the LLMs’ output-format. Specifying the output format improves the efficiency and accuracy of the results extraction in certain tasks.
The bolded font is used as the name of the modules in the introduction list.
3.3 Internal Basic Elements
Three purposes are typically included in prompts: (1) Implying a certain message to LLMs; (2) Letting LLMs execute a certain task with or without output; (3) The combination of the first two.
[…]
Module Samples of Basic Elements
Profile
$\bullet$ You are a magazine editor.
Goal
$\bullet$ You need to generate a title for the article.
Constraint
$\bullet$ The length of the title should not exceed 20 words.
Workflow
### Extracting the kernel content
[…]
$\circ$ Analyse the theme of the article;
$\circ$ Detecting the main objects and related things described in the article;
$\circ$ Summarising the core content from the article;
$\circ$ Save the kernel content.
Style
$\bullet$ The style of the title should be formal.
Table 2: Examples of basic internal elements of inherent modules in the writing scenario. This prompt belongs to the writing category and its purpose is to generate a title for a given article. We have chosen five modules as examples - profile, goal, constraint, workflow, and style - and show one internal element from each module. Particularly, for the workflow module, we show a function-like basic element.
Figure 2: Design process for extension modules and custom elements. For application scenarios where inherent modules cannot cover all aspects, new extension modules can be defined as needed. After defining the extension module, the internal elements should also be designed according to the requirements of the extension module. In addition, if inherent modules are compatible with the needs of the scenario but the basic elements cannot meet all of the requirements, custom elements can be added directly to the inherent modules. It is important to note that extension and inherent modules should be as mutually exclusive as possible to minimize modification costs.
3.4 Extension Module and Custom Element
The inherent modules we have defined have been as comprehensive as possible to cover the many aspects of prompts. Furthermore, we add application scenarios and modules covered by LangGPT. However, limited by our own capabilities and domain knowledge, we were unable to consider all application scenarios of LLMs.
Therefore, in addition to inherent modules and basic elements, we define the extension module and custom element to improve the generalization and reusability of prompts. The design flow of extension modules and custom elements is shown in Figure 2.
This process helps users quickly analyze their needs and come up with the key points that need to be included in the prompt. Based on the design templates provided by LangGPT and this design process, quality prompts can be efficiently designed to match the application scenarios.
4 Experiments
To validate the advancement of the proposed LangGPT, we conducted experiments in two aspects: the tasks executed by LLMs and the usability of LangGPT. In our experiments, we chose two types of application scenarios: writing and role-playing. For each scenario, we selected five task-specific assistants constructed and shared by users in our community.
Figure 3: Ratings on ease of use in user survey. The lowest score is 0, which means very difficult to use, and the highest score is 5, which means very easy to use. The “:” is used to separate scores and percentages.
4.1 Baselines
For comparison, we chose two baselines.
  • Instruction-only. The prompts contain only the instructions proposed for LLMs and necessary information as needed.
  • CRISPE prompt design rules Nigh (2023). The framework prospectively defines the components of the prompt from a macro perspective. It requires a complete prompt containing Capacity and Role, Insight, Statement, Personality, and Experiment.
Instruction-only prompts are very intuitive and maximize token savings. However, this approach contains too little content, often fails to convey the full range of requirements, and makes it difficult to focus on key points. CRISPE provides a relatively complete design framework rather than a number of fragmented rules. However, it does not have a clear structured design, which is not conducive to learning and the reuse of high-quality prompts. In addition, CRISPE has a strict specification of the elements to be included in the prompts, having a low generalization capability.
Figure 4: A case of a boot-licker. The responses of ChatGPT-3.5 to the user under three different prompts. Mingyuan University doesn’t really exist.
Because of the shortcomings of these methods in prompt design, we designed LangGPT. However, in addition to these intuitive advantages, we need to verify the ability of prompts to induce LLMs, which is the most essential purpose of prompt design.
4.2 Large Language Models
We selected these LLMs***All executive tasks in the experiment were completed by 18 January 2024. shown in Table 3 for evaluation.
LLM Access Scale Version
ChatGPT Schulman et al. (2023) API unk v3.5
ChatGLM Du et al. (2021) API unk v3-turbo
ChatGLM Zeng et al. (2023) Open Source 6B v3
Ernie Bot Sun et al. (2021) API unk v4.0
Gemini Anil et al. (2023) API unk Pro
Baichuan Yang et al. (2023) Open Source 7B v2-chat
Qwen Bai et al. (2023) Open Source 7B chat
Yi 01-ai (2023) Open Source 6B chat
Table 3: LLMs used in experiments. “unk” indicates that we do not know the size of those models.
4.3 Evaluation Metrics
Due to the complexity of the task scenarios, there is a lack of objective evaluation metrics. Therefore, we evaluated the ability of LLMs to execute tasks using human evaluation and LLM evaluation. To better ensure the rationality of the evaluation, we set evaluation criteria for two scenarios.
For writing tasks, we asked the evaluators to assess the effectiveness of the LLMs in performing the task in terms of three dimensions: textual continuity, formatting norms, and content richness. The ability to generate thematically coherent content is important. However, we found that LLMs in the generation task were usually able to maintain thematic consistency, so no evaluation metrics were designed for this aspect. For role-playing tasks, we similarly designed 3 dimensions of evaluation metrics: language style, characteristic relevance, and thematic coherence.
For each evaluation metric, we defined the score from zero to five to indicate different levels and indicated the description of the situation corresponding to each integer score. In addition, we allowed evaluators to score 0.5 between levels.
This evaluation framework can also guide high-performance LLMs, such as GPT-4 and Ernie Bot-4, to automatically evaluate their performance. As a result of the low consistency and of LLMs as evaluators Liu et al. (2023), we evaluate them manually.
4.4 Results
The most intuitive manifestation of the ability of a prompt to induce LLMs is the performance of LLMs in performing tasks guided by the prompt. Thus, we evaluated the performance of LLMs in two scenarios and the results are shown in Table 4.
Scenario Prompt GPT-3.5 GLM-turbo ChatGLM3 Ernie Bot Gemini Pro Baichuan2 Qwen Yi
s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3 s1 s2 s3
Writing I 3.8 3.5 3.7 3.5 3.1 3.2 3.7 3.5 3.8 4.2 3.9 3.7 3.8 4.2 4.5 3.5 3.6 3.3 4.1 3.7 4.4 4.1 3.8 4.3
C 3.7 3.1 3.4 3.5 3.1 3.1 3.9 3.7 4.1 4 3.9 4 4.1 4.1 4.1 3.7 3.6 3.4 4.3 4 4.2 4.3 4 4.2
L 3.8 3.5 3.4 3.4 3.4 3.4 3.5 4 3.8 4.4 4 4.2 4.2 3.9 4.1 3.6 3.4 3.7 3.4 3.6 3.7 3.9 4.1 4.6
Role-playing I 4.2 3.5 3.4 3.8 1.6 1.4 3.4 1.4 1.7 4 0.8 2.1 4.4 2.8 3.4 3.9 1.3 1.2 3.9 2.9 2.8 3.6 2 2.4
C 4 3.7 4 4.2 2.4 2.5 4 3 3.7 4.3 2.6 2.5 4.6 4.2 4 4.1 2.5 2.4 4.3 3.2 3.4 4.2 3.2 3
L 4.3 4.3 4.3 4.2 1.8 2.1 4.1 4 4.3 4.2 2.3 2.5 4.6 4.4 4.3 4.4 3 3.2 4.2 3.7 3.7 3.8 2.8 2.8
Table 4: Results of different prompts inducing LLMs to perform tasks. Bolded fonts indicate the type of prompts that perform best in the current scenario. I, C, and L denote three kinds of prompts, i.e., Instruction-only, CRISPE, and LangGPT, respectively. $\text{s}_{1}$ to $\text{s}_{3}$ represent the three evaluation metrics for the two scenarios, respectively. For the writing scenario, $\text{s}_{1}$ is textual continuity, $\text{s}_{2}$ is formatting norms, and $\text{s}_{3}$ is content richness. For the role-playing scenario, $\text{s}_{1}$ is language style, $\text{s}_{2}$ is characteristic relevance, and $\text{s}_{3}$ is thematic coherence.
Scenario Writing Role-playing
Prompt I C L I C L
GPT-3.5 3.67 3.40 3.57 3.70 3.90 4.30
GLM-turbo 3.27 3.23 3.40 2.27 3.03 2.70
ChatGLM3 3.67 3.90 3.77 2.17 3.57 4.13
Ernie Bot 3.93 3.97 4.20 2.30 3.13 3.00
Gemini Pro 4.17 4.10 4.07 3.53 4.27 4.43
Baichuan2 3.47 3.57 3.57 2.13 3.00 3.53
Qwen 4.07 4.17 3.57 3.20 3.63 3.87
Yi 4.07 4.17 4.20 2.67 3.47 3.13
Table 5: The average of the scores of the different LLMs on the two types of Scenarios.
From the results, it can be seen that LangGPT is better at inducing LLMs to execute tasks. In addition, we noticed an interesting phenomenon. Certain LLMs with particularly strict security restrictions will refuse to answer questions or requests on topics such as bragging and crankiness. LangGPT can make these LLMs realize that it’s just a style of speaking rather than truly damaging behavior and guide them to respond. Neither of the other two baselines can do that.
To evaluate the ease of use of LangGPT, we conducted a user survey in our online communityThe community has been running for more than six months and has amassed thousands of users from a wide range of industries, including manufacturing, construction, information technology, finance, and entertainment. Therefore, the objectivity of the survey results can be guaranteed.. We designed a complete questionnaire about the LangGPT experience to ensure the quality of answers. The questionnaire included a rating question on ease of use. The results of the user ratings are shown in Figure 3.
As can be seen from Figure 3, 87.81% of users gave a score of 3 or higher, which indicates users’ approval of LangGPT’s ease of use. In addition, LangGPT’s overall satisfaction score in the user survey was 8.48 out of 10.
4.5 Case Study
To demonstrate the effect of LangGPT more intuitively, we filtered specific cases from our experiments. In addition to the direct effect comparison, we also tried to use LangGPT to guide LLMs to generate quality prompts.
4.5.1 Inducing LLMs
In our experiments, we guided the LLMs to play a boot-licker using three prompts, an example of which is given in Figure 4.
In this example, both the Instruction-only prompt and the CRISPE prompt-guided ChatGPT simply reply to the user’s utterances. In contrast, LangGPT-guided ChatGPT is even more bottomless in its blow-by-blow approach to the user-given subject. Additionally, it expresses compliments from a wider range of perspectives and is more in character.
Figure 5: Example of ChatGPT-3.5 to generate prompts with LangGPT. Some modules have been omitted for ease of presentation.
4.5.2 Prompt construction with LLMs
In organizing our experiments, we also tried to make LLMs construct prompts using LangGPT. An example is shown in Figure 5.
If LLMs are directly asked to generate a prompt about MBTI assessment, they might reject it. However, LangGPT can direct LLMs to generate high-quality prompts that are harmless.
5 Conclusion
In this paper, we present LangGPT, a structured and extensible framework for prompt design. LangGPT has a systematic structure similar to object-oriented programming languages and is easy to learn and reuse. Experiments demonstrate that LangGPT performs better than the baseline methods in guiding LLMs to perform tasks. We also conducted a user survey in the community built on LangGPT to verify the ease of use and reusability of LangGPT. In future work, we will further optimize the design of LangGPT and reduce the token consumption of LangGPT. In addition, support for LLMs using third-party tools and custom tools will be added.
Ethical Statement
In the application of LLM, ethical disputes may arise, but the design of LangGPT and the process of writing this paper avoided possible ethical issues.